OpenStudy (anonymous):

I'm stuck on a homework questions. Please help! In the Bohr model, an electron in a hydrogen atom makes the following transitions. Which of the following would produce the light with the shortest wavelength in the emission spectrum? n = 1 to n = 5 n = 3 to n = 2 n = 6 to n = 5 n = 4 to n = 3

OpenStudy (anonymous):

You have to use Roberts equation right?

OpenStudy (anonymous):

Sorry Rybergs

OpenStudy (anonymous):

I don't know rybergs equation yet

OpenStudy (anonymous):

There's a energy level diagram that can help

OpenStudy (anonymous):

where can i find it?

OpenStudy (anonymous):

also there was another one that i'm stuck on.....how can i calculate how far does light travel in....1.5secs, 9.0 days and 1.0 yr? what is the formula to figure that out?

OpenStudy (anonymous):

Rydberg equation is \[ \frac{1}{\lambda} = R \Bigg( \frac{1}{n_1^2} - \frac{1}{n_2^2} \Bigg) \] \[n_1<n_2\] Also, non-analytically, transitions at lower levels yield higher energy photons.

OpenStudy (anonymous):

as for the second question, light in a vacuum travels at \[c=3 \times 10^8 \ m/s\] So convert 1.5secs, 9.0 days and 1.0 yr into seconds, and the formula to find distance, d, is \[vt = d \quad \overset{\text{ in this case}}{\longrightarrow} \quad ct=d\]

OpenStudy (anonymous):

thank you!!!